Summary of the GMVA Session, March 31 - April 5, 2033 TPK; 13.4.2022 ***************************************************** Stations participating: Eb, On, Mh, Ys, Pv, Nn, GLT VLBA(8) @3mm, VLBA(10 @7mm GBT @3mm 3xKVN + MOPRA ATCA @ 1Gbps ALMA maser just broke before the observations started, so there is ono ALMA in this session. LMT withdraw on short notice due to lack of man-power and security issues in personal transportation Haystack joined to for some hours for testing ====================================================================================== European Fringetest: -------------------- f221a - Fringes found to Eb, On, Mh Nn, and later during the session: Ys, PV Overall Summary: ----------------- In Europe the session was parly affected by snow and wind at the high sites (IRAM), rain and wind at Effelsberg, and bad weather in the 2nd half of the session at Mh and On. The 2nd half ofd the session was under good weather conditons at Noema. Weather at VLBA was good, except BR and NL. LA had a warm receiver, PT had electronic problems. GBT missed the first few hours on c221a (ML010A, M87) due to internet problems. They also were not recording for 16.75 hrs on c221c (MG009, OJ287). GLT was obsering but had some period of snow stops. =============================================================================================== Detailed Notes from Stations: Effelsberg: ----------- Effelsberg was affected by relatively poor weather (rain, storm) in the beginning and at the end of the session. Technically the observations went well, but the telescope could not observe because of high wind on several occasions, as listed below: f221a - start 2 scans late, rest is ok c221a - weather degrading after Apr.1, 3 UT, no pointing possible after Apr.1, 13 UT c221b - telescope stowed because of snow, resume observing Apr.2, 10 UT with scan No103 c221c - start on time, weather ok c221d - start on time, on Apr.4, 14 UT telescope stopped due to wind (Scan No0070) until end of c221d c221e - start om time, poor weather after Apr. 5, 3 UT GLT: ----- notes from Slack, #c221 channel Apr. 1 0 UT GLT begins, snow 3 UT stop, snow 14 UT resume obs. Apr. 2 17 UT light snow Apr. 4 8 UT GLT good, Tsys~230K at elv=20, tau=0.11 23 UT GLT begins c221e Note after correlation: GLT uses a wrong LO tuning in part B,C and D, no fringes ! Metsahovi: ---------- A short summary (Mh) about session. - No technical issues, expect minor issue in fringe test (f221a). Chopper motor was freezing and we missed scans 02 and 03. - f221a, c221a, c221b and c221c were under good weather conditions. - c221d and c221e were under snow storm, especially c221e was under awful weather. - No missed scans on c221a - c221e. All the best, Juha PS. I will put log- and antabfs-files to the server within next days. NOEMA: ------- notes from Slack, #c221 channel March 31: 16 UT stop, high wind 2- UT resume obs., clear sky Apr. 1 : 7 UT stop, snow 16 UT stop, snow 22 UT stop, ice on dish Apr. 2 12 UT stop, bad weather Apr. 3 20h30 UT good, Tsys=70-150K Apr. 5 07 UT sessions ends with good weather Onsala ------- Onsala has successfully participated in the GMVA session 1/2022. As you know, there were high-SNR fringes to the Onsala 20-m telescope with the correct polarisation in the first ftp-fringe test of F221A. Five scientific experiments were done in the extremely good weather: blue sky, dry air and Tsys* 130 K at EL >~70 deg. Only the experiment C221D suffered the rain (till 16 UT) and had Tsys* >~ 300 K. All the antabfs files have been uploaded to VLBEER. Before the session, the pointing mode was updated by our local expert Henrik Olofsson. I did a few pointing observations. It looked good (error <10 arcsec). The phase should be stablely locked through the whole 3mm session. I did not get any error email from our antenna control software. At Onsala, we had a spike, provided by a spectrometer, to monitor the phase stability. It was always clearly seen. After the session, the test signal was checked again to verify the LO correctness and the phase stability accurately. Regards, Jun Yang Pico Veleta: -------------- notes from Slack, #c221 channel: March 31: 15 UT bad weather, no obs. Apr. 1 6 UT still bad, no obs. 11h30 UT PV joins c221a 16 UT clear sky, Tsys~140 K Apr. 2 7 UT good weather, Tsys~ 130K, tau=0.08 Apr. 3 18 UT good weather, Tsys~ 140K, tau=0.12 Apr. 4 8 UT good weather, Tsys~ 140K, tau=0.08 13 UT Pico stops, snow 20h30 UT Pico stops, wind+snow Apr. 5 02h30 UT resume obs., Tsys~160K fog+clouds Here is a summary of the GMVA Spring Session 2022 at Pico Veleta. The weather during the session was a mix of bad, and good + excellent weather. For the individual schedules: f221a: Lost completely due to bad weather (Snow, Wind) c221a: A large fraction lost due to weather (Snow, Wind). Storm then passes and sky becomes clear, with low opacity tau225~0.15. Start recording at 12:00UTC. Done 11.3% of the schedule. c221b: Done 100% under excellent weather c221c: Done 100% with some fog and thin clouds, not affecting much the quality of the data. c221d: Done first part with cloudy sky. Have to stop in scan no 0177 due to snow and wind. Done 66.3% of the schedule. c221e: Bad weather keeps antenna stowed except for a window of ~1.5 hr when we can record with fog and some clouds. Done 23.7% of the schedule. You can look up the exact scans done/lost in vlbi@mrt-enter.iram.es:~/vlbireduc/mar2022/3mm/vlbi_mar2022_3mm_log The calibration results and byproducts (like ANTAB tables) will be produced later. We will let you know when everything is available. With best regards, The Pv VLBI team (Pablo, Salvador, Ignacio) Yebes: ------ technical problems in the beginning, first half of c221a probably bad, fringes found Apr.1, 08 UT Yebes 40m RT used linear polarization until Friday 1, April at 16:00. After 16:00 we inserted a quarter wavelength plate and the telescope started using circular polarization. We did pointing and focus during some gaps on SiO masers close to the target sources and applied corrections. KVN: ---- Here is a feedback of KVN sites. Sky conditions during this session were good. There were some problems at Ky and Ku as follows. Ky: 093 05:30 - 093 07:30 large offset in sub-reflector position due to operational mistake 094 12:40 - 095 00:51 stop due to antenna driver failure Ku: 092 13:40 - 092 14:04 field system problem 095 16:34 - 095 17:30 field system problem Kt : no problem I've uploaded station logs and antab files on the evn server. ATCA: ------ Due to an error ATCA recorded as USB/USB, not LSB/USB. Note the cover sky frequency is still correct, we just forgot to 'flip' the USB setup for the lower frequency to LSB. This flip occurs at the centre frequency of each band individually, so does not effect the recorded sky frequency. This is 100% recoverable in DIFX - just change the LSB or USB in the $FREQ section and subtract 64 MHz from the effective LO freq (band edge) also in the $FREQ section. GBT: ----- internet problems in the 1st experiment (Apr 1, 0-7 UT), start observations a few hours later (ML010A in c221a) Logs and Tsys are here: https://safe.nrao.edu/wiki/bin/view/GB/Observing/WbandVLBACal -> C221A to C221D VLBA, summary per experimentcode: ===================================== PT and LA have problems LA 3mm is warm VERY LONG BASELINE ARRAY OBSERVING LOG -------------------------------------- Project: C221A Observer: Krichbaum, T. Project type: OTHER Obs filename: c221a.vex Date/Day: 2022MAR31/090 Ants Scheduled: SC HN NL FD LA PT KP OV BR MK GB =UT-Time===Comment===============================================MF#===%AD==AMD= Operator is Betty Ragan 1653 Begn at all stes except SC,HN,GB 1653-1800 *LA 3mm cryo warm. (WO-13088) CRYO 50 33 1653-1728 *PT M401 and L404 are having communiations MCB 100 35 problems after DBE replacement. Site techs are working on it. 1659 Begin at SC,HN 1728-1755 *PT M401 is OK now, but the L404 keeps SYN 50 13 crashing. Things might at least be set up correctly for 7mm scans. Tsys for those are high, but not unreasonable. 1755 PT has everything up again for the moment, although some 3mm tsys look implausibly low, so there might still be a problem. It is also likely the L404 will crash again. There seems to be some sort of networking problem. 1800-1919 *LA parked for site techs to work on 3mm cryo. CRYO 100 79 (WO-13088) 1818-2122 *PT autostowed. Looks like it had a bad power COM 100 184 hit. 1822 PT removed from observation while we investigate what just happened. 1912 PT reports commercial power lost a phase. No ETA on when it will be restored. Continuing to leave it out. 1919-2352 *LA back in. 3mm is still warm. (WO-13088) CRYO 50 136 2122-1345 *PT power issues have been resolved, but now SYN 100 983 the L404 is completely dead. (WO-13095) 2120-2140 *HN FRM having trouble getting to correct FRM 50 10 position. Keeps having drive faults and ending up ~0.3 degrees or so off. 2210 New operator: Jessica King 2350 GB begin. 2350-0112 *GB files are failing to transfer. Executor cannot SW 100 82 load project. 0000 Date is 2022APR01/091. 0040 %HN precipitation. 0317 %SC precipitation. 0532 New Operator: Paul Padilla 0655 %MK raining 1150-1301 *HN multiple FRM errors. Not clearing with a FRM 100 71 Master Clear. Contacting the Site Techs. 1153 GB ends 1157 HN parked and out of observation until Techs can get on site to investigate. 1319-1321 *KP autostowed, power glitch? COM 100 2 1345 PT L404 apparently fixed itself. Putting it back into observation. 1345 New operator is Betty Ragan 1345-1355 *PT tsys garbage at 3mm on the first SYN 50 5 few scans. 1345-2352 *PT 7mm cryo warm. (WO-13111) CRYO 50 303 1533 %LA raining intermittently 1537-1557 *LA parked for site techs to work on cryo. CRYO 100 20 1602-1617 *PT parked for site techs to work on L404. SYN 100 15 (WO-13095) 1613-1622 *LA parked for site techs again. CRYO 100 9 1804-1821 *LA parked for site techs again. CRYO 100 17 1806-1952 *HN,MK out for USNO observation. SCH 200 212 2010 End at SC 2155 Operator: Jessica King 2154-2352 *PT L404 has died again. Switches are not being set SYN 50 59 correctly for 3mm scans. 2352 End. Downtime Summary: Total downtime : 2268 min Percentage downtime of observing: 12.2% Average downtime per hour : 7.3 min Total scheduled observing time (# Antennas): 18590 min (10) VERY LONG BASELINE ARRAY OBSERVING LOG -------------------------------------- Project: C221B Observer: Krichbaum, T. Project type: OTHER Obs filename: c221b.vex Date/Day: 2022APR01/091 Ants Scheduled: SC HN NL LA PT KP OV BR MK GB =UT-Time===Comment===============================================MF#===%AD==AMD= Oeprator: Jessica King 2250 GB begin. 2354 NL,FD,LA,PT,KP,OV,BR begin. 2354-0007 *PT L404 failure. Switches are not being set SYN 50 6 correctly for 3mm. 2354-2353 *LA 3mm warm. CRYO 50 719 0000 Date is 2022APR02/092. 0007-0105 *PT removed for L404 work. SYN 100 58 0105-2353 *PT L404 was replaced and is communicating. FE 50 684 Switches are being set correctly. However, 3mm data is still bad. Tsys are high, and RMS counts are around 2.5. At this point I suspect there may be an issue with the amplification process, but we can't be sure without further inspection. 0354 Begin at MK 0545 New Operator: Paul Padilla 0653-0655 *PT autostowed, power glitch? COM 100 2 0722-0834 *MK out for scheduled Q-series USNO observation SCH 100 72 0853-1511 %NL precipitation. 1345 New operator is Alan Kerr 1545 Begin at SC,HN. 1800-1816 *LA out to work on a cryo. CRYO 100 16 2349 End at SC. 2350 End at GB. 2353 End. Downtime Summary: Total downtime : 1557 min Percentage downtime of observing: 11.5% Average downtime per hour : 6.9 min Total scheduled observing time (# Antennas): 13527 min (9) VERY LONG BASELINE ARRAY OBSERVING LOG -------------------------------------- Project: C221C Observer: Krichbaum, T. Project type: VLBA Obs filename: c221c.vex Date/Day: 2022APR02/092 Ants Scheduled: SC HN NL FD LA PT KP OV BR MK GB =UT-Time===Comment===============================================MF#===%AD==AMD= Operator is Alan Kerr Full Observation began at 1900 2354 Begin at NL,FD,LA,PT,KP,OV,BR 2354-2354 *LA 3mm warm (96K). CRYO 50 720 0000 Date is 2022APR03/093. 0007 Begin at GB. 0007-1643 *GB intermittent communication. not recording. SW 100 996 looks similar to Xcube problems from C221A. 0109 Begin at MK. 0145 New operator is Carolyn Morris 0722-0831 *MK out for USNO SCH 100 69 1016-1019 *MK autostow. power glitch? COM 100 3 1340 New operator is Alan Kerr 1552-1609 *LA out to work on a cryo. CRYO 100 17 1721-1819 *LA out to work on the cryo again. CRYO 100 58 1807-1951 *HN,MK out for scheduled USNO observation. SCH 200 208 2254 End at SC,HN 2331 End at GB 2354 End. Downtime Summary: Total downtime : 2071 min Percentage downtime of observing: 14.4% Average downtime per hour : 8.6 min Total scheduled observing time (# Antennas): 14400 min (10) VERY LONG BASELINE ARRAY OBSERVING LOG -------------------------------------- Project: C221D Observer: Krichbaum, T. Project type: VLBA Obs filename: c221d.vex Date/Day: 2022APR04/094 Ants Scheduled: SC HN NL FD LA PT KP OV BR MK GB =UT-Time===Comment===============================================MF#===%AD==AMD= Operator is Alan Kerr Observation began at 0830 on Apr03 0009 Begin at NL,FD,LA,PT,KP,OV,BR 0009 %NL raining 0009-0054 *LA 3mm warm (194K) CRYO 100 1485 0015 Begin at SC,HN 0041 Begin at MK. 0200 %MK raining 0200 New operator is Carolyn Morris 0752 Begin at GB 1318 End at GB 1345 New operator is Alan Kerr 2200 New operator is Betty Ragan 2313 %OV windy 0000 Date is 2022APR05/095. 0054 End. Downtime Summary: Total downtime : 1485 min Percentage downtime of observing: 10.0% Average downtime per hour : 6.0 min Total scheduled observing time (# Antennas): 14850 min (10) VERY LONG BASELINE ARRAY OBSERVING LOG -------------------------------------- Project: C221E Observer: Krichbaum, T. Project type: VLBA Obs filename: c221d.vex Date/Day: 2022APR05/095 Ants Scheduled: NL FD LA PT KP OV BR MK GB =UT-Time===Comment===============================================MF#===%AD==AMD= Operator is Betty Ragan 0051 Begin at GB 0053 Begin at NL 0104 Begin at FD,LA,PT 0104-1229 *LA 3mm cryo warm. (WO-13088) CRYO 50 342 0104-0115 %FD is complaining that record rate is zero, but none of the other sites that started at the same time is recording yet, so I think it isn't actually supposed to be, either. 0124 Begin at KP 0158 Begin at OV,BR 0426 Begin at MK 0435 End at GB 0533-0535 *OV autostowed. Power glitch. COM 100 2 0600 New operator: Jessica King 1113 NL end. 1212 FD end. 1229 End. Downtime Summary: Total downtime : 344 min Percentage downtime of observing: 6.2% Average downtime per hour : 3.7 min Total scheduled observing time (# Antennas): 5584 min (8) Notes: * = Entries where data was affected. % = Entries where data may have been affected. & = Entries where the site tech was called out. WEA = Weather entries. MF# = Maintenance form or major downtime category associated with a problem. %AD = The percentage of an antenna affected by a problem. AMD = Total antenna-minutes downtime for a problem. Tsys = System Temperature (TP/SP x Tcal/2) ACU = Antenna Control Unit FRM = Focus/Rotation Mount RFI = Radio Frequency Interference VME = Site control computer CB = Circuit Breaker vclock = Program that compares site clock time to a standard.